Complete the squre for the parabola y=3x^2+5x+17
Answer by ewatrrr(24785)

 
Hi,
Complete the square for the parabola
y=3x^2+5x+17 = 3(x+5/3)^2 - 25/3 + 17 = 3(x+5/3)^2 + 26/3